OpenStudy (anonymous):

what is the equation of a line with a slope of -7 and a y-intercept of 6

OpenStudy (anonymous):

any ideas

Vocaloid (vocaloid):

m = slope b = y-intercept your equation: y = mx + b plug in your slope and y-intercept

OpenStudy (anonymous):

?

Vocaloid (vocaloid):



Vocaloid (vocaloid):



Vocaloid (vocaloid):

now plug in the numbers you are given for slope and y-intercept

Vocaloid (vocaloid):

replace "m" with your slope replace "b" with your y-intercept

OpenStudy (anonymous):

a.6y=-7x b.y=6x+(-7) c.-7y=x+6 d.y=-7+6 would it be c?

Vocaloid (vocaloid):

no, please do what I asked

OpenStudy (anonymous):

i know i had an answer before